# Spectral Appearance of Self-gravitating AGN Disks Powered by Stellar Objects: Universal Effective Temperature in the Optical Continuum and Application to Little Red Dots

### v1

@@ I Introduction
Since its launch, the James Webb Space Telescope (JWST) has uncovered a previously unrecognized class of compact, red sources known as “Little Red Dots” (LRDs; Matthee et al. 2024), which are now being identified in large numbers. These objects are unresolved in rest-optical images (e.g., Kokorev et al., 2024; Akins et al., 2025) and show broad Balmer emission lines (e.g., Greene et al., 2024; Lin et al., 2024) together with a distinctive V-shaped continuum: they are blue in the rest-frame ultraviolet (UV) but extremely red in the rest-frame optical (Labbé et al., 2023; Kocevski et al., 2024; Hviding et al., 2025), with the red optical continuum dominating the total spectral output (Greene et al., 2025). Their extreme compactness and unusual spectral energy distributions have attracted considerable attention, motivating a range of new physical scenarios for their nature and evolution.
At first glance, the most straightforward interpretation is a conventional dust-reddened AGN accretion disk plus a stellar population component from the host galaxy (e.g. Wang et al., 2024; Labbe et al., 2024; Ma et al., 2025a; Wang et al., 2025a). However, the substantial dust extinction implied by these models is difficult to reconcile with the stringent constraints on dust emission in the rest-mid and far-infrared (FIR) from deep JWST/MIRI and ALMA observations (Williams et al., 2024; Akins et al., 2025; Casey et al., 2025; Setton et al., 2025; Wang et al., 2025a; Xiao et al., 2025). This motivates scenarios in which the optical/red bump arises from thermal emission produced by dense, gas-dominated envelopes in an almost dust-free environment. Such configurations resemble magnified versions of stellar envelopes at their Hayashi limits with a characteristic $T_{\rm eff}\sim 5000$K regulated by H- opacity, and have been referred to as “black hole stars” or “quasi-stellar”-type models (de Graaff et al., 2025a; Naidu et al., 2025; Begelman & Dexter, 2025; Kido et al., 2025; Liu et al., 2025; Nandal & Loeb, 2025; Wang et al., 2025c). This class of models is also motivated by strong Balmer breaks observed in a subset of LRDs (Inayoshi & Maiolino, 2025; Naidu et al., 2025; Ji et al., 2025; de Graaff et al., 2025b), a feature that poses a serious challenge for both standard AGN continua and conventional post-starburst stellar populations. On the blue side, the lack of strong rest-UV continuum variability in most systems (Furtak et al., 2025; Ji et al., 2025; Zhang et al., 2025b, c) adds further tension with a conventional AGN picture and provides important additional constraints on viable physical models.
@@ III.2 Integration Procedure
and within $R_{\rm sg}$ the disk should connect to a standard disk (where we expect no star formation to further reduce the mass inflow) with a constant $\dot{M}(R_{\rm sg})$ and an integrated luminosity of $\sim\epsilon_{\bullet}\dot{M}(R_{\rm sg})c^{2}$ in the UV/X-ray, with the radiative efficiency $\epsilon_{\bullet}$ determined by the standard AGN inner cutoff close to a few gravitational radii $R_{g}$. In our picture of LRDs, this emission should be dominated by outer stellar emission regardless of the inner disk structure (e.g. see Figure 8). But for completeness, we do solve a standard disk structure within $R_{\rm sg}$ using the viscous disk equations (Shakura & Sunyaev, 1973) assuming a constant accretion rate.
@@ IV.1 Fiducial Models at $M_{\bullet}=10^{6}M_{\odot}$ and Implied Constraints
Eventually at sufficiently large $\gamma$, a qualitative transition occurs: the disk reaches an optically thin branch before the self-gravitating radius, at a radius $R_{\rm thin}>R_{\rm sg}$. This behavior is evident in the $\gamma=0.8$ case for $\dot{M}=0.1M_{\odot}$/yr in Figure 3 (green opaque lines). This is due to the strong depletion of $\dot{M}$ that yields $\partial T(R)/\partial R>0$ within a certain radius and makes $T_{\rm eff}$ and $T$ converge inwards. When this happens, we terminate the iterative solution at $R_{\rm thin}$. We do not attempt to model the detailed physics in this optically thin region which is uncertain, same as the reason we do not model the optically thin cloud exterior to $R_{\rm out}$.
@@ VI.3 Cosmological Abundances
where $t_{\rm LRD}$ is the characteristic duration of the accretion phase (that can be observed), ${\dot{M}_{\rm acc}}/{\dot{M}_{\rm lum}}$ denotes the ratio between the total mass accreted and the mass converted into radiation, which is $(1-\epsilon_{\bullet})/\epsilon_{\bullet}\sim 1/\epsilon_{\bullet}$ for accretion-powered source. Multi-band observations of LRDs have constrained their bolometric luminosities to be substantially lower than earlier estimates based on standard AGN bolometric corrections (Greene et al., 2025). With these recently revised luminosities, adopting a timescale of $t_{\rm LRD}\sim 10^{9}{\rm years}$ corresponding to the cosmic timescale at $z\sim 5$, and a canonical radiative efficiency $\epsilon_{\bullet}\sim 0.1$, the implied mass density $\rho_{\rm LRD}\lesssim 10^{4}-10^{5}M_{\odot}$Mpc-3 for $\mathcal{L}_{\rm LRD}\sim 10^{40}-10^{41}$erg s-1 Mpc-3 would be below the local SMBH mass density $\rho_{\rm BH}\sim 3\times 10^{5}M_{\odot}$Mpc-3 estimated from other methods (e.g. Bernardi et al., 2010, 2013; Liepold & Ma, 2024) and does not pose a significant Sołtan-type tension. In fact, it would be consistent with the interpretation that LRDs generally transition into AGNs around $z\sim 5$, a redshift below which their abundance undergoes a significant decline (Ma et al., 2025b; Umeda et al., 2025).

### v2

@@ I Introduction
Since its launch, the James Webb Space Telescope (JWST) has uncovered a previously unrecognized class of compact, red sources known as “Little Red Dots” (LRDs; Matthee et al. 2024), which are now being identified in large numbers. These objects are unresolved in rest-optical images (e.g., Kokorev et al., 2024; Akins et al., 2025; Hviding et al., 2025; de Graaff et al., 2025a) and show broad Balmer emission lines (e.g., Greene et al., 2024; Lin et al., 2024) together with a distinctive V-shaped continuum: they are blue in the rest-frame ultraviolet (UV) but extremely red in the rest-frame optical (Labbé et al., 2023; Kocevski et al., 2024; Hviding et al., 2025), with the red optical continuum dominating the total spectral output (Greene et al., 2025). Their extreme compactness and unusual spectral energy distributions have attracted considerable attention, motivating a range of new physical scenarios for their nature and evolution.
At first glance, the most straightforward interpretation is a conventional dust-reddened AGN accretion disk plus a stellar population component from the host galaxy (e.g. Wang et al., 2024; Labbe et al., 2024; Ma et al., 2025a; Wang et al., 2025a). However, the substantial dust extinction implied by these models is difficult to reconcile with the stringent constraints on dust emission in the rest-mid and far-infrared (FIR) from deep JWST/MIRI and ALMA observations (Williams et al., 2024; Akins et al., 2025; Casey et al., 2025; Setton et al., 2025; Wang et al., 2025a; Xiao et al., 2025). This motivates scenarios in which the optical/red bump arises from thermal emission produced by dense, gas-dominated envelopes in an almost dust-free environment. Such configurations resemble scaled-up versions of stellar envelopes at their Hayashi limits with a characteristic $T_{\rm eff}\sim 5000$K regulated by H- opacity, and have been referred to as “black hole stars” or “quasi-stellar”-type models (de Graaff et al., 2025a; Naidu et al., 2025; Begelman & Dexter, 2025; Kido et al., 2025; Liu et al., 2025; Nandal & Loeb, 2025; Wang et al., 2025c). This class of models is also motivated by strong Balmer breaks observed in a subset of LRDs (Inayoshi & Maiolino, 2025; Naidu et al., 2025; Ji et al., 2025; de Graaff et al., 2025b), a feature that poses a serious challenge for both standard AGN continua and conventional post-starburst stellar populations. On the blue side, the lack of strong rest-UV continuum variability in most systems (Furtak et al., 2025; Ji et al., 2025; Zhang et al., 2025b, c) adds further tension with a conventional AGN picture and provides important additional constraints on viable physical models.
@@ III.2 Integration Procedure
and within $R_{\rm sg}$ the disk should connect to a standard disk (where we expect no star formation to further reduce the mass inflow) with a constant $\dot{M}(R_{\rm sg})$ and an integrated luminosity of $\sim\epsilon_{\bullet}\dot{M}(R_{\rm sg})c^{2}$ in the UV/X-ray, with the radiative efficiency $\epsilon_{\bullet}$ determined by the standard AGN inner cutoff close to a few gravitational radii $R_{g}$. In our picture of LRDs, this emission is dominated by outer stellar emission regardless of the inner disk structure (e.g. see Figure 8). But for completeness, we do solve a standard disk structure within $R_{\rm sg}$ using the viscous disk equations (Shakura & Sunyaev, 1973) assuming a constant accretion rate.
@@ IV.1 Fiducial Models at $M_{\bullet}=10^{6}M_{\odot}$ and Implied Constraints
Eventually at sufficiently large $\gamma$, a qualitative transition occurs: the disk reaches an optically thin branch before the self-gravitating radius, at a radius $R_{\rm thin}>R_{\rm sg}$. This behavior is evident in the $\gamma=0.8$ case for $\dot{M}=0.1M_{\odot}$/yr in Figure 3 (green opaque lines). This is due to the strong depletion of $\dot{M}$ that yields $\partial T(R)/\partial R>0$ within a certain radius and makes $T_{\rm eff}$ and $T$ converge inwards. When this happens, we terminate the iterative solution at $R_{\rm thin}$. We do not attempt to model the detailed physics in this optically thin region which is uncertain, just as we do not model the optically thin cloud exterior to $R_{\rm out}$.
@@ VI.3 Cosmological Abundances
where $t_{\rm LRD}$ is the characteristic cosmic time over which LRDs are abundantly observed, ${\dot{M}_{\rm acc}}/{\dot{M}_{\rm lum}}$ denotes the ratio between the total mass accreted and the mass converted into radiation, which is $(1-\epsilon_{\bullet})/\epsilon_{\bullet}\sim 1/\epsilon_{\bullet}$ for accretion-powered source. Multi-band observations of LRDs have constrained their bolometric luminosities to be substantially lower than earlier estimates based on standard AGN bolometric corrections (Greene et al., 2025). With these recently revised luminosities, adopting a timescale of $t_{\rm LRD}\sim 10^{9}{\rm years}$ corresponding to the cosmic timescale at $z\sim 5$, and a canonical radiative efficiency $\epsilon_{\bullet}\sim 0.1$, the implied mass density $\rho_{\rm LRD}\lesssim 10^{4}-10^{5}M_{\odot}$Mpc-3 for $\mathcal{L}_{\rm LRD}\sim 10^{40}-10^{41}$erg s-1 Mpc-3 would be below the local SMBH mass density $\rho_{\rm BH}\sim 3\times 10^{5}M_{\odot}$Mpc-3 estimated from other methods (e.g. Bernardi et al., 2010, 2013; Liepold & Ma, 2024) and does not pose a significant Sołtan-type tension. In fact, it would be consistent with the interpretation that LRDs generally transition into AGNs around $z\sim 5$, a redshift below which their abundance undergoes a significant decline (Ma et al., 2025b; Umeda et al., 2025).